NASA is conducting an experiment to find out the fraction of people who black out at G forces greater than 6.

Step 2 of 2 :  

Suppose a sample of 543543 people is drawn. Of these people, 217 passed out. Using the data, construct the 98% confidence interval for the population proportion of people who black out at G forces greater than 6. Round your answers to three decimal places.

phat = 0.400
Answer #1

Confidence Intervals: Confidence Intervals are one way to draw the conclusion about the population based on a random sample data.

  • The Population Proportion is denoted by 'p'.
  • The Sample Proportion is denoted by '\hat{p} '. (Natural Estimate)
  • The Sample Size is denoted by 'n'.

Constructing an Interval around \hat{p} which has higher probability including p, is known as Confidence Interval (CI).

To arrive at the confidence interval, we have to know Estimate \pm Margin of Error which is formulated as

  \hat{p}\pm z* SE(\hat{p})

In an Interval, It is expressed as (\hat{p} - z* SE(\hat{p}) , \hat{p} + z* SE(\hat{p}))

The factor z* is based on the normal distribution.

  • \hat{p} = 0.400
  • p = 217/ 543543 = 0.000399
  • SE = sqrt [ p(1 - p) / n ] = sqrt [0.000399(1-0.000399)/ 543543] = 7.315
  • z*=2.33 from t-distribution calculation

Therefore, The confidence interval range is 0.4 \pm2.33(7.315) = 0.4 \pm17.044

The confidence interval is 16.644 to 17.444